### v2

# Comparing Hyper-optimized Machine Learning Models for Predicting Efficiency Degradation in Organic Solar Cells

Abstract
This work presents a set of optimal machine learning (ML) models to represent the temporal degradation suffered by the power conversion efficiency (PCE) of polymeric organic solar cells (OSCs) with a multilayer structure ITO/PEDOT:PSS/P3HT:PCBM/
Al. To that aim, we generated a database with 996 entries, which includes up to 7 variables regarding both the manufacturing process and environmental conditions for more than 180 days. Then, we relied on a software framework that brings together a conglomeration of automated ML protocols that execute sequentially against our database by simply command-line interface. This easily permits hyper-optimizing and randomizing seeds of the ML models through exhaustive benchmarking so that optimal models are obtained. The accuracy achieved reaches values of the coefficient determination (R2) widely exceeding 0.90, whereas the root mean squared error (RMSE), sum of squared error (SSE), and mean absolute error (MAE) $\lesssim$1of the target value, the PCE. Additionally, we contribute with validated models able to screen the behavior of OSCs never seen in the database. In that case, R2$\sim$0.96-0.97 and RMSEsimilar-$\sim$1%, thus confirming the reliability of the proposal to predict. For comparative purposes, classical Bayesian regression fitting based on non-linear mean squares (LMS) are also presented, which only perform sufficiently for univariate cases of single OSCs. Hence they fail to outperform the breadth of the capabilities shown by the ML models. Finally, thanks to the standardized results offered by the ML framework, we study the dependencies between the variables of the dataset and their implications for the optimal performance and stability of the OSCs. Reproducibility is ensured by a standardized report altogether with the dataset, which are publicly available at Github [^1] https://github.com/davidumh11.
keywords:
Organic solar cell, Degradation, P3HT:PCBM, PEDOT:PSS, Machine Learning, Regression.
journal: Engineering Applications of Artificial Intelligence
[…]
[label1]organization=University Institute for Engineering Research, Miguel Hernandez University,addressline=Avenida de la Universidad, s/n, city=Elche, postcode=03202, country=Spain
[…]
[label2]organization=Communications Engineering Department, Miguel Hernandez University,addressline=Avenida de la Universidad, s/n, city=Elche, postcode=03202, country=Spain
[…]
[label3]organization=Institute of Chemical Synthesis and Homogeneous Catalysis (ISQCH), CSIC, University of Zaragoza,addressline=Pedro Cerbuna 12, city=Zaragoza, postcode=50009, country=Spain
1 Introduction
Nowadays electricity consumption has evolved exponentially over the last three decades, increasing from a total estimate of 15107.8 TWh to 28661 TWh between 2000 and 2022, respectively Rit2023 . One of the immediate consequences has been reflected in the price of this energy. Simultaneously, in the technological realm, there has been a notable shift in electricity generation towards renewable sources. Particularly, the maturity of conventional photovoltaic (PV) cells and the emergence of new organic and hybrid materials have made possible to get significant improvements in PV generation and its power conversion efficiency (PCE) NREL . However, despite the fact that the global consumption of PV energy increases from 1.08 TWh to 1310.02 TWh between 2000 and 2022 Rit2023 , its generation still falls short of meeting current demand. Hence, there are many challenges and open lines for improving the generation and efficiency of PV cells.
In this context, organic solar cells (OSCs) have emerged as a promising alternative to silicon-based solar cells since, amongst other benefits, their production involves low-temperature manufacturing methods Seo2015 , reducing their carbon footprint. Their appeal also lies in the ease of processing, low cost, flexibility, and lightweight Yeh2013 . Moreover, the evolution of their PCEs has been much superior to classical technologies and it has been on par with other considered emerging technologies, presenting efficiencies that exceed values of 20% veintePCE . However, the commercialization of OSCs still faces various challenges, including stability and efficiency. Nevertheless, certain cost studies PVCost1 ; PVCost2 indicate that once these obstacles are overcome, OSCs can be manufactured at a cost lower than 1 dollar per peak Watt Bra2005 . This fact suggests that the optimization of the manufacturing procedure of these organic devices would certainly lead to reinforce their benefits, but also, and most importantly, their sustainability and stability to operate long-term feasibly. All in all, the potential capabilities of artificial intelligence (AI) protocols, through particular machine learning (ML)-based models, envisages great achievements in this sense.
The integration of ML before transferring to industrial implementation it is being extensively tested these days. For instance, within the field of cheminformatics, it is of paramount importance in the design of new materials. Numerous studies have demonstrated the potential of deep learning algorithms to analyze complex chemical data, thereby accelerating the process of discovering, for example, drugs, as well as identifying promising compounds with improved properties Lus2013 . Some of the early studies Han2013 ; Rup2012 laid the foundations for the application of ML models in predicting molecular properties, highlighting their great strengths. More recently, the deployment of AI techniques with high-performance prediction methods has enabled the emergence of software platforms capable of examining extensive chemical databases to identify potential materials with significant properties Unt2014 ; Wall2015 . A relevant database, which moved forward from general chemical structures to the inclusion of PV features was the project presented in Joh2011 . This work allowed to obtain custom predictive models that characterize the electrical behavior of PV devices in terms of parameters such as voltage, current, conversion efficiency, short-circuit current, open-circuit voltage, etc. Eib2021 ; Ser2016 ; Tol2019 ; Lop2016 . Thus there is an endless list of mathematical models mostly supported by ML and deep learning-based engines that combine different techniques, such as regression and statistical inference Moo2022 ; Eib2021 ; Mal2021 ; Seo2019 ; Sun2019 to optimize different target features of optoelectronic devices Miy2021 ; Mah2022a ; Mah2022b . Similarly, the growth of proposals with eco-design characteristics and AI support for the production of sustainable and durable energy generating materials is positively evolving Vel2023 ; Ju2018 ; Mah2023 . More specifically, some studies have trusted on ML for predicting optimal installation of energy harvesting systems HeliyonReview1 ; PV_forecast_EAAI ; HeliyonReview2 and consequently on predicting energy generation HeliyonPV ; ML_predictUSA ; PV_forecast_EAAI . Similarly, fault detection of these systems were also anticipated thanks to automated models DefectVisual ; DefectEAAI . Others delved into the manufacturing variables of their specific structures ML_structure ; Power_predict ; HeliyonPerov ; ML_OSC . In the same way, there is incipient research on the analysis of stability and durability of these devices PerovBig ; Lifetime ; PEDOT_stability ; PerovML .
In sight of all these developments, this work explores different ML models to encode the PCE performance of polymeric OSCs based on multilayer ITO/PEDOT:PSS/P3HT:PCBM/Al. This permits producing models that reliable learn from the PCE evolution over time. Furthermore, several dependencies on the variables involved in the manufacturing PhysicaB_Dopant_study , but also on the environmental conditions DegradationVariables ; ThermalAnalisis ; DegradationHumidity , are also learnt by these models. To that aim, we have used an hyper-optimizer software, ROBERT ROBERT , that permits benchmarking with multiple ML models and so that to produce extensive and standardized results, which are generated with optimal methods. Our dataset consists of PCE measurements of 45 OSC devices with variability in up to 7 variables, acquired for more than 180 days. This means a database with 996 entries. The accuracy metrics demonstrate the validity of these models to learn and represent the temporal behavior of our OSCs, with determination coefficient up to R2=0.96, and root mean squared error (RMSE), sum of squared error (SSE) and mean absolute error (MAE) bounded around 1% of the target value, that is, the PCE. Additionally, traditional Bayesian regression models sustained by LMS have been introduced in order to confirm the advantages of the ML-automated protocols against these classical approaches. While these conventional methodologies such as PVBayes ; PVStatistical ; PVStatistical2 may yield satisfactory results in this domain, our approach demonstrates that the potential of computational learning modeling offers a much broader and comprehensive solution compared to those based on classical statistical models, showcasing, among other benefits, increased robustness and reliability for:
  • Capturing multivariate relationships between variables, encompassing both linear and non-linear relationships.
  • Learning from an entire dataset, not solely from a specific OSC device as in classical regression.
  • Screening new OSC devices not used during the learning phase, thereby predicting its behavior, as an unknown device.
  • Predicting optimal variables to fabricate the OSC that maximizes the PCE and/or its stability over time.
In contrast ML models, typical statistical algorithms are unable to predict; instead, they utilize estimations to extrapolate target values to immediately subsequent time instances, which must also be exclusive to a particular OSC rather than to an entire database with variability across different variables.
In summary, the key contributions of this research are as follows:
  1. 1.
    Assessment of optimal ML models that learn the PCE behavior over time for polymer-based OSC devices with ITO/PEDOT:PSS/P3HT:PCBM/Al structure, considering various fabrication variables and environmental measurement conditions.
  2. 2.
    Comparative benchmarking among different ML models and classical statistical regression approaches.
  3. 3.
    Identification of an optimal ML model capable of accurately predicting the PCE behavior of unseen OSC devices, unknown during the training and validation phases.
  4. 4.
    Analysis of dependencies among the dataset variables to establish influence on the performance of the OSCs devices and their fabrication.
  5. 5.
    Reproducibility and transparency of the obtained ML models using a standardized framework for command line replication.
The rest of the paper is organized as follows: Section 2 describes the specific OSCs manufactured in our laboratory, their electrical parameters, and the periodic measurements that comprise the database used by the ML methods. Subsequently, these methods are defined, starting from a preliminary scope of regression problems, along with the software framework that enables their benchmarking and, consequently, the extraction of models with optimal hyper-parameters. Then, Section 3 outlines the experiments and results. Finally, Section 4 draws conclusions from this work.
2 Materials and Methods
2.1 Organic Solar Cells
The OSCs characterized in this work are manufactured through the spin-coating technique, which involves the deposition of overlapped polymeric thin films. Polymers are deposited in solution, then a process of rotation followed by heating aids in removing the solvent, hence forming the layer. The structure of the devices was as follows: ITO/PEDOT:PSS/P3HT:PCBM/Al, where each layer refers to:
  • ITO: Indium tin oxide.
  • PEDOT:PSS: poly(3,4-ethylenedioxythiophene) polystyrene sulfonate.
  • P3HT:PCBM: (poli(3-hexiltiofeno-2,5-diil):[6,6]-phenyl-C61-butyric acid methyl ester.
  • Al: Aluminium.
As a substrate, a glass with a 60 nm thick semi-transparent ITO layer was used. The substrates are placed in the spinner with the ITO layer facing upwards so that the thin layers can be deposited on it. The first layer deposited was the PEDOT:PSS film. PEDOT:PSS was deposited at room temperature at a rotation speed of 6000 rpm for 60 seconds. The remaining solvent was removed by heating at 150C for ten minutes. Once the first layer was deposited, the active layer composed by the P3HT:PCBM polymer blend was applied. This was done at 300 rpm for three minutes. It was then dried at 80C for one hour.
The final layer deposited on the devices is the aluminium film. Aluminium is not deposited using the spin-coating technique but rather by metal evaporation in a vacuum chamber. The equipment consists of a high vacuum chamber, two vacuum pumps, one of which is a high vacuum pump, and a power supply to provide the necessary current to evaporate the aluminium. Once the samples are placed inside the chamber, vacuum is achieved, reaching pressures of 10-6 mbar. Once this pressure is reached, aluminium is evaporated by Joule effect.
2.2 OSC database
As initially mentioned in Section 1, our dataset consists of PCE measurements of 45 OSC devices where up to 7 variables are registered for more than 180 days. Electrical parameters (J-V curves) of these OSCs are acquired along with climate conditions. The acquisition process comprises a Keithley-2400 equipment that acts as source generator for the voltage sweep as well as recorder of the generated PV current. The J-V curves have to be measured under light conditions (100 mW/cm2, AM 1.5G, and 25C) generated by a solar simulator Newport xenon arc lamp and an AM 1.5G filter. The electrical characterization is completed by determining the characteristic electrical parameters, including the short-circuit current density (Jsc), the open-circuit voltage (Voc), the maximum power point (Pmpp), the fill factor (FF), and the PCE. Obtaining these parameters allows to extract the PCE value as follows:
$PCE=\frac{P_{mpp}}{P_{inc}}=\frac{J_{sc}V_{oc}FF}{G}$ (1)
being Pinc the incident solar power on the OSCs, which derives from the incident irradiance of the solar simulator, G.
The block diagram of the equipment for the acquisition system is presented by Figure 1. This permitted obtaining a database with 996 entries. The variables and their units are presented by Table 1.
Figure 1: Block diagram of the acquisition system used to generate the OSC database.
Table 1: Variable details of the dataset.
Variables Values
Cell name [1-5]
solvent quantity HTL (PEDOT:PSS) [${\mu}$l] [250-1000]
P3HT [mg] [1-1.2]
PCBM [mg] [0.8-1]
Volume ratio P3HT:PCBM [1-1.25]
Temperature [C] [12-23]
Hummidity [%] [33-88]
Dew point [C] [3-19]
Pressure [hPa] [997-1022]
Time [days] [0-181]
As an example, Figure 2 presents an OSC contained in the database, for which its current density (J) data have been periodically acquired against voltage (V) over 78 days. Specifically, Figure 2 displays the voltage range on the X-axis from 0 to 0.5 V, while Figure 2 adjusts the scale to show the range from 0 to 0.25 V, allowing for a clearer observation of the evolution of the J-V curve as the days pass by. The degradation principles of these organic devices suggest that over time, these curves, from which maximum power and energy conversion efficiency are obtained, should consistently decline. However, it can be observed in Figure 2 that there are certain temporal points between days 20, 26, 47 and 69, where the current density associated with specific voltages remains close to, or even exceeds, the current density associated with those same voltages but at earlier temporal points. This outline confirms that rather than time, there are other variables influencing the degradation of these devices, which implies that a specific OSC could exhibit slightly better performance in the future under certain conditions.
Figure 2: Evolution of current density (J) vs. voltage (V) over time (days) for the OSC $Cell2$. (a) Curve J-V with V $\in$[0-0.7] V. (b) Same curve J-V with V $\in$[0-0.25] V.
In the same manner, Figure 3 compares the temporal evolution of three different OSCs in terms of their normalized PCE values, ranging from 0 to 1, for more than 180 days. Once again, it is confirmed that time leads to the degradation of the device’s conversion efficiency. However, it also emerges that differences in the fabrication of the various OSCs, as well as the environmental conditions during their measuring, may influence the trend followed by the PCE over time.
Figure 3: Evolution of the normalized PCE over time for three different OSCs.
2.3 Regression Problem
As already stated, this study aims to automatically model the degradative behavior exhibited by manufactured OSCs using ML approaches. It is clear that we deal with a regression problem, for which recent literature has demonstrated that is reliably and efficiently solved by computational learning. However, in order to establish further comparisons that highlight and reinforce the abilities of AI-based protocols, a preliminary step is carried out to present regression fitting using traditional LMS-supported Bayesian methods.
Regression is a method for estimating the relationship between a response or output variable and one or more predictor or input variables. Linear and non-linear regression serve as estimators of values between observed data points. From that starting point, a regression model relates response data to predictor data through one or more coefficients. Then a fitting algorithm is needed to calculate some model coefficients from a set of input data. In this sense, parametric fitting involves finding those coefficients, assuming that the data are statistical composed by deterministic and random components. Therefore, the parametric model estimates the deterministic component, whereas the random component is typically described as the error. Considering the model as a coefficient function of the independent variable, the error encodes random fluctuations around a Gaussian probability distribution. In many scopes, the goal is to minimize that error, classically addressed by means of LMS fitting approaches, such as: linear least squares, weighted least squares, robust least squares, non-linear least squares, etc.
$y=f(X,\beta)+\varepsilon$ (2)
where $y$ is the output vector data of n×$\times$1, corresponding to the input data in $X$ of n×$\times$×m, after being applied $f$ as a non-linear function of the coefficient vector $\beta$ (m×$\times$1), being $\varepsilon$ the vector of unknown errors of n×$\times$1. Afterwords, the SSE is minimized, understood as the residual sum of squares, given a set of $n$ data values, the residual value of the $i$-th value $r_{i}$ is calculated as:
$r_{i}=y_{i}-\hat{y}_{i}$ (3)
where $y_{i}$ represents the $i$-th observed value and $\hat{y}_{i}$ represents the $i$-th estimated value, and accordingly:
$SSE=\sum_{i=1}^{n}r_{i}^{2}=\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (4)
Subsequently, the algorithm proceeds iteratively calculating the coefficients from an initial seed. Sometimes non-linear models trust on heuristic schemes to calculate initial values. For others models, coefficients randomly initialized in ranges from [0-1]. Then the response value is given as $\hat{y}=f(X,\beta)$, computed using the jacobian matrix of $f(X,\beta)$, as the matrix that contains the partial derivatives with respect to the coefficients of $\beta$. Finally, the adjustment of the coefficients for the next iteration lies on some non-linear least squares algorithms, such as Levenberg-Marquardt, Gradient descent or Gauss-Newton LMS . Whenever the fitting meets the specified convergence criteria, the final solution is assumed as valid.
After observing in Section 2.2 the behavior of PCE over time for the various OSCs contained in our database, it was validated that the LMS regression fitting models yielding the best results were those with non-linear characterization. Table 2 displays the selected models along with their expressions as a function of time, dependent on the adjustment coefficients. It should be noted that the capability of these classical models lies solely in modeling the univariate behavior of the time effect on the PCE values. While it is possible to explore other LMS regression fitting in the multivariate domain, they only allow for establishing linear relationships, which do not adequately capture the behavior of our devices.
Table 2: LMS Bayesian regression fitting models to estimate PCE, denoted as $f(x)$, where $x$ represents time by means of parametric models.
Parametric model Coefficient expression of $f(x)$
$exp1$ $ae^{bx}$
$exp2$ $ae^{bx}+ce^{dx}$
$gauss1$ $a_{1}e^{-[(x-b_{1})/c_{1}]^{2}}$
$gauss2$ $a_{1}e^{-[(x-b_{1})/c_{1}]^{2}}+a_{2}e^{-[(x-b_{2})/c_{2}]^{2}}$
$poly3$ $p_{1}x^{3}+p_{2}x^{2}+p_{3}x+p_{4}$
2.4 Machine Learning framework
In contrast to the previous classical approaches, ML moves forward to produce non-parametric regression models that adjust more complex behaviors without specifying the relationship between the output and input data by certain predetermined regression function. They may simply predict responses with new test data using a trained model. In this work, we exploit the advantages of a software framework, ROBERT ROBERT , developed under Python that facilitates hyper-optimization and benchmarking over well-recognized ML regression models by single command line instruction. This framework consists of the following modules:
  • Data curation: It processes the target dataset in order to filter correlated descriptors, noise, duplicates, as well as to identify and to convert categorical variables into one-hot descriptors.
  • Model selection: It iterates through multiple hyper-optimized models from scikit-learn pedregosa11a . It includes training size analysis by varying data partitions percentages (training-validation) and it also examines different data splitting options like k-neighbors clustering-based (KN) and random (RND). Permutation feature importance (PFI) detection is also performed. In consequence, models are computed with PFI-filtered descriptors and without them. As a default setting, this threshold eliminates features whose contribution to the model’s R2 is less than 4%. Moreover, it provides with insights on the dependencies of the input variables, thanks to SHapley Additive exPlanations (SHAP) analysis. Besides this, heatmap graphics are presented in order to compare models in terms of RMSE error. Amongst the generated models, it assesses the performance of Random Forests (RF), Multivariate Linear Models (MVL), Gradient Boosting (GB), Gaussian Process (GP), AdaBoost Regressor (AdaB), MLP Regressor Neural Network (NN), and Voting Regressor (VR) pedregosa11a .
  • Prediction: Once the models are trained and validated, the framework permits external test sets to predict target values with the best models.
  • Verify module: It assesses the predictive ability of the models, considering tests such as y-shuffle, y-mean, k-fold cross-validation, and one-hot features prediction.
  • Report module: With the aim of providing reproducibility and transparency, this module offers a detailed report containing comprehensive information about the ML models utilized and replication instructions through command line executions.
As an example of performance, Figure 4 illustrates the result obtained by ROBERT with our dataset as input data. Specifically, Figure 4 and Figure 4 compare the evolution exhibited by the ML models over time. The first figure depicts the result of a single model for training data up to 30 days, while the second figure does the same for training data up to 180 days. Their validity can be checked by inspecting the predicted outcome against the target values used as training (blue) and validation (yellow) data.
Figure 4: Results of the best ML models, tested with training (blue) and validation (yellow) target data up to: (a) 30 days and (b) 180 days.
3 Results
This section introduces the results obtained using both classical LMS regression fitting models and the optimal ML models as generated by the framework used in this work, to estimate the temporal behavior of our OSCs. The selected error metrics for analysis are briefly presented below:
  • Coefficient of determination R2: It quantifies how well the independent variables explain the variability of the dependent variable. Higher values indicate that the model fits the data well and captures larger proportion of the variability in the dependent variable.
    $R^{2}=1-\frac{SSE}{SS_{tot}}$ (5)
  • RMSE: It provides a measure of the average magnitude of the errors made by the model in its predictions. Minimizing this error is often a goal when training regression models in ML.
    $RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$ (6)
  • MAE: Another common objective error that measures the average magnitude of the errors between the predicted values and the actual values of the target variable. Unlike RMSE, which penalizes large errors more heavily, MAE treats all errors equally by taking the average of their absolute values. This makes MAE more robust to outliers in the data compared to RMSE.
$MAE=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$ (7)
Meanwhile, SSE was defined in (4) and SStot is the total sum of squares, whereas $n$ is the number of samples, $y_{i}$ is the real observed value and $\hat{y}_{i}$ is the predicted value.
3.1 Classical LMS regression fitting results
Hence, Figure 5 presents the accuracy metrics of the five Bayesian regression models introduced in Table 2 to estimate the performance of PCE over 180 days, for each of the OSCs available in the database (Table 1). It is worth noting that these results correspond to the mean values, along with the standard deviation for each OSC. Specifically, the values of R2 are shown in Figure 5, RMSE in Figure 5, SSE in Figure 5, and MAE in Figure 5. It can be observed that these models perform effectively when the temporal range of the fitting data is limited: as long as the time goes by, all models experience an increase in error, while R2 decreases. According to the results of Figure 5, it is observed that on average, the $gauss2$ model, consisting of two Gaussian terms, is the one that best estimates the performance of the OSCs in terms of PCE over time. However, it should be noted that for a 30 day fit, this model requires more points to achieve a valid R2. On the other hand, it should be pointed out that the maximum errors for all temporal fittings are bounded to values $\sim$0.06, fact that represents an error $>$1% of the the target value, the PCE. Notwithstanding that, it must be reiterated the limitation of these methods to characterize the behavior of all devices contained in our database in a global manner. These results reflect solely the mean performance of fitting for single OSCs, in turn.
Figure 5: Accuracy metrics of the LMS-supported Bayesian regression fitting over time. Five different models are evaluated: $exp1$blacksquare, $exp2$blacksquare, $gauss1$blacksquare, $gauss2$blacksquare, $poly3$blacksquare. (a) Coefficient of determination, R2. (b) Root mean squared error, RMSE. (c) Sum of squared errors, SSE. (d) Mean absolute error, MAE.
Figure 6: PCE estimation over time with the best LMS-supported Bayesian regression fitting model ($gauss2$). Four temporal datasets are used to compute the fittings: 30 days $\circ$-- ∘ -; 60 days $\circ$-- ∘ -; 90 days $\circ$-- ∘ - and 120 days $\circ$-- ∘ -.
3.1.1 PCE fitting estimation
Once preliminary fitting have been made with traditional LMS regression methods, it has been found that the best fitting corresponds to a double-term Gaussian function. Now we analyze the robustness of this method to estimate the behavior of PCE against the temporal variable. For this purpose, different estimations have been obtained, as shown in Figure 6. This figure presents estimated PCE data from a specific OSC device, namely $Cell4$ in our databased, which has been used so far in the all the previous examples throughout this work. Here, the parametric model is obtained with fitting data at: 30, 60, 90, and 120 days, respectively. Then the behavior of PCE at 180 days has been estimated, in order to be tested against the real data of the database. It can be observed that the estimated trend for fitting models with 30, 60, and 90 days deviate considerably from the real behavior (represented by the black line). It can be concluded that fitting data up to at least 120 days are required to make a robust fitting that achieves sufficient confidence in estimating future data at 180 days. This fact demonstrates that even the best parametric LMS fitting regression model requires more than half of the temporal data to reliably model the PCE evolution at future time values. In order to get further insights on the estimations plotted in Figure 6 more easily, Table 3 presents the specific details of these estimations according to the error metrics RMSE, SSE, and MAE. Upon inspection, the perception extracted from the previous figure is validated: the errors are unacceptable for any metric unless the fit is made with more than half of the data (90 days), and only if the subsequent PCE estimation is tested over time values immediately posterior to the temporal limit of the fitting.
Table 3: Accuracy metrics of the best LMS-supported Bayesian regression fitting model ($gauss2$) to estimate PCE over time, presented in Fig 6.
Fitting data Estimation RMSE SSE MAE
30 days 60 days 0.0516 0.0213 0.0277
90 days 0.1141 0.1302 0.0680
120 days 0.1872 0.4905 0.1339
150 days 0.2224 0.8405 0.1705
180 days 0.2442 1.3117 0.1979
60 days 90 days 0.0454 0.0206 0.0220
120 days 0.1083 0.1641 0.0685
150 days 0.1511 0.3879 0.1046
180 days 0.1911 0.8038 0.1445
90 days 120 days 0.0150 0.0032 0.0103
150 days 0.0160 0.0043 0.0119
180 days 0.0190 0.0079 0.0118
120 days 150 days 0.0158 0.0042 0.0127
180 days 0.0138 0.0038 0.0105
3.2 Machine Learning regression results
In this section, we conduct the generation of ML models to learn from the temporal behavior of PCE versus time. It is noteworthy that, unlike the previous LMS-based regression fittings presented in Section 3.1, these ML models allow us to characterize the performance of our devices in a multivariate manner, thereby encoding the variability presented in their manufacturing, as well as in the particular environmental conditions at each temporal measurement, as indicated in Table 2.
Thanks to the use of the tool ROBERT ROBERT , we achieve an exploration of various combinations of ML algorithms and partition sizes. In order to maintain the same accuracy estimation policy used in Section 3.1, Figure 7 presents the same metrics as Figure 5. In this case, for each temporal training dataset (X-axis), the coefficient R2 is presented in Figure 7, RMSE in Figure 7, SSE in Figure 7, and MAE in Figure 7, for the best ML model in each case. It can be observed that, in general, all errors are bounded approximately within ranges $\sim$[0.02-0.03], for training data of 120 days onwards. This implies an error percentage lower than 1%, with respect to the mean value of PCE for these OSCs, which is highly satisfactory for modeling. It is also worth noting the good performance of the RF model, with training-validation partitions of 90-10%.
So as to comparatively assess the remaining ML models, Tables 45678, and 9, respectively present detailed accuracy results for the other models (GB, NN, MVL, and RF). Results with different training-validation ratios are presented: from 90-10% to 60-40%; while training data range from 30 to 180 days. For each table, the robustest model is highlighted in bold, considering the overall performance across all presented metrics, with priority given to the R2 value. It is worth mentioning that, even though they are not directly comparable models, it is confirmed that the errors of the ML methods are clearly bounded, while the LMS-based fitting regression errors grow with time. Furthermore, ML models have the ability to globally model a multivariate database, in contrast to univariate Bayesian fitting, which is independent for each device.
Figure 7: Accuracy of the best Machine Learning regression models over time: 30 days RF-60-10 $\blacksquare$; 60 days RF-90-10 $\blacksquare$; 90 days RF-90-10 $\blacksquare$; 120 days RF-90-10 $\blacksquare$; 150 days RF-90-10 $\blacksquare$ and 180 days RB-90-10 $\blacksquare$. (a) Coefficient of determination, R2. (b) Root mean squared error, RMSE. (c) Sum of squared errors, SSE. (d) Mean absolute error, MAE.
Table 4: Accuracy metrics of the Machine Learning regression models computed with training data up to 30 days.
Training data ML model R2 RMSE SSE MAE
30 days GB-80-20 0.86 0.140 0.0647 0.1200
GB-70-30 0.73 0.140 0.0622 0.1200
GB-60-40 0.84 0.160 0.1103 0.1500
NN-80-20 0.76 0.180 0.1655 0.1800
NN-70-30 0.51 0.170 0.0832 0.1600
NN-60-40 0.62 0.170 0.0931 0.1400
MVL-80-20 0.71 0.200 0.0984 0.1500
MVL-70-30 0.62 0.220 0.1006 0.1400
MVL-60-40 0.66 0.170 0.1067 0.1400
RF-80-20 0.86 0.130 0.0573 0.1000
RF-70-30 0.78 0.130 0.0558 0.1000
RF-60-40 0.89 0.120 0.0725 0.0950
Table 5: Accuracy metrics of the Machine Learning regression models computed with training data up to 60 days.
Training data ML model R2 RMSE SSE MAE
60 days GB-90-10 0.85 0.1000 0.0155 0.0800
GB-80-20 0.76 0.1500 0.0433 0.1300
GB-70-30 0.84 0.1500 0.0597 0.1300
GB-60-40 0.78 0.1700 0.1275 0.1500
NN-90-10 0.89 0.1300 0.1239 0.1000
NN-80-20 0.86 0.1700 0.1677 0.1500
NN-70-30 0.88 0.1500 0.0506 0.1300
NN-60-40 0.79 0.1600 0.0677 0.1400
MVL-90-10 0.71 0.1500 0.1059 0.1300
MVL-80-20 0.80 0.1700 0.1120 0.1500
MVL-70-30 0.63 0.1700 0.1117 0.1500
MVL-60-40 0.61 0.1900 0.1290 0.1500
RF-90-10 0.98 0.1000 0.0702 0.0850
RF-80-20 0.92 0.1400 0.0832 0.1300
RF-70-30 0.78 0.1600 0.0747 0.1400
RF-60-40 0.81 0.1300 0.0578 0.1000
Table 6: Accuracy metrics of the Machine Learning regression models computed with training data up to 90 days.
Training data ML model R2 RMSE SSE MAE
90 days GB-90-10 0.87 0.0730 0.0255 0.0660
GB-80-20 0.87 0.1100 0.0736 0.0860
GB-70-30 0.88 0.0940 0.0775 0.0770
GB-60-40 0.75 0.1400 0.0755 0.1100
NN-90-10 0.86 0.0790 0.0332 0.0630
NN-80-20 0.82 0.1300 0.1595 0.1200
NN-70-30 0.46 0.1900 0.1071 0.1700
NN-60-40 0.52 0.1900 0.1611 0.1500
MVL-90-10 0.61 0.1100 0.3161 0.1000
MVL-80-20 0.63 0.1700 0.1468 0.1400
MVL-70-30 0.54 0.1700 0.1378 0.1400
MVL-60-40 0.55 0.1800 0.1569 0.1500
RF-90-10 0.89 0.0750 0.0300 0.0670
RF-80-20 0.87 0.0940 0.0909 0.0700
RF-70-30 0.80 0.1000 0.0613 0.0780
RF-60-40 0.74 0.1400 0.0860 0.1100
Table 7: Accuracy metrics of the Machine Learning regression models computed with training data up to 120 days.
Training data ML model R2 RMSE SSE MAE
120 days GB-90-10 0.95 0.0670 0.0312 0.0540
GB-80-20 0.88 0.0700 0.0324 0.0510
GB-70-30 0.89 0.0810 0.0453 0.0620
GB-60-40 0.81 0.1300 0.1447 0.1100
NN-90-10 0.88 0.0600 0.2140 0.0450
NN-80-20 0.83 0.0890 0.0576 0.0660
NN-70-30 0.71 0.1300 0.0767 0.0960
NN-60-40 0.59 0.1700 0.1743 0.1300
MVL-90-10 0.70 0.0990 0.1729 0.0850
MVL-80-20 0.65 0.1400 0.1820 0.1100
MVL-70-30 0.60 0.1600 0.2311 0.1400
MVL-60-40 0.51 0.1800 0.2370 0.1500
RF-90-10 0.98 0.0300 0.0507 0.0240
RF-80-20 0.88 0.0680 0.0651 0.0480
RF-70-30 0.86 0.0920 0.0609 0.0690
RF-60-40 0.84 0.1000 0.0574 0.0770
Table 8: Accuracy metrics of the Machine Learning regression models computed with training data up to 150 days.
Training data ML model R2 RMSE SSE MAE
150 days GB-90-10 0.98 0.0300 0.0510 0.0230
GB-80-20 0.93 0.0660 0.0242 0.0420
GB-70-30 0.94 0.0730 0.0351 0.0580
GB-60-40 0.91 0.0820 0.0591 0.0640
NN-90-10 0.90 0.0580 0.0760 0.0540
NN-80-20 0.82 0.0110 0.1277 0.0630
NN-70-30 0.88 0.0960 0.0871 0.0740
NN-60-40 0.78 0.1200 0.1533 0.0910
MVL-90-10 0.84 0.0920 0.2113 0.0800
MVL-80-20 0.87 0.1100 0.2139 0.0840
MVL-70-30 0.88 0.1200 0.2163 0.0900
MVL-60-40 0.82 0.1100 0.2184 0.0910
RF-90-10 0.97 0.0290 0.0219 0.0210
RF-80-20 0.96 0.0510 0.0394 0.0324
RF-70-30 0.94 0.0780 0.0417 0.0570
RF-60-40 0.90 0.0930 0.1069 0.0943
Table 9: Accuracy metrics of the Machine Learning regression models computed with training data up to 180 days.
Training data ML model R2 RMSE SSE MAE
180 days GB-90-10 0.96 0.0640 0.0595 0.0520
GB-80-20 0.96 0.0720 0.0600 0.0560
GB-70-30 0.94 0.0840 0.0672 0.0650
GB-60-40 0.95 0.0700 0.0431 0.0500
NN-90-10 0.66 0.1400 0.3732 0.1100
NN-80-20 0.81 0.1200 0.1067 0.0710
NN-70-30 0.87 0.0860 0.0824 0.0660
NN-60-40 0.89 0.0990 0.1021 0.0800
MVL-90-10 0.81 0.1000 0.2929 0.0810
MVL-80-20 0.80 0.1100 0.2645 0.0890
MVL-70-30 0.78 0.1200 0.2626 0.0870
MVL-60-40 0.74 0.1300 0.2802 0.0920
RF-90-10 0.96 0.0280 0.0246 0.0290
RF-80-20 0.94 0.0690 0.0391 0.0510
RF-70-30 0.94 0.0670 0.0398 0.0460
RF-60-40 0.93 0.0850 0.0644 0.0580
3.2.1 PCE prediction
After comparatively evaluating the performance of ML models, this section validates their ability to predict the temporal behavior of the PCE for an OSC device never seen by the models. In other words, a predictive inference of PCE is obtained for data from a particular OSC that has been omitted in the training and validation phases. Firstly, the models are trained and validated without data of $Cell4$ (same example used so far), thus the accuracy metrics are presented in Table 10. Standing out above the others are the GB-90-10 and RF-90-10 models. Secondly, the four trained models presented in Table 10, are queried to get predictive inferences up to 180 days for the unseen OSC, i.e. $Cell4$. Table 11 summarizes these accuracy results for each of these inferences. Again, both in Table 10 (training-validation) and in Table 11 (predictive testing) the highlighted error metrics together with the R2 values higher than 0.90, are indicative of feasibility to achieve our goal. Finally, Figure 8 provides visual detail for the particular results obtained with one of the best models, GB-90-10, to predict PCE over 180 days. Since the figure only shows the behavior of PCE against the time variable, once again, it is worth emphasizing on the potential of these models to encode accurately the multivariate variations of our dataset for an unknown OSC device.
Table 10: Accuracy metrics of the Machine Learning regression models without training nor validation data of the OSC $Cell4$, which is not seen by the models.
ML model R2 RMSE SSE MAE
GB-90-10 0.88 0.069 0.0906 0.0600
NN-90-10 0.89 0.095 0.3005 0.0740
MVL-90-10 0.80 0.091 0.7528 0.0790
RF-90-10 0.96 0.071 0.1118 0.050
Table 11: Accuracy metrics of the Machine Learning regression models used as predictors of the PCE for the unseen OSC, $Cell4$.
ML model R2 RMSE SSE MAE
GB-90-10 0.97 0.055 0.0914 0.0370
NN-90-10 0.80 0.1000 0.1975 0.0800
MVL-90-10 0.90 0.1500 0.4010 0.1500
RF-90-10 0.94 0.0610 0.0674 0.0470
Figure 8: PCE prediction over time with the best Machine Learning regression model (GB-90-10, Table 11) for an OSC not seen during the training neither the validation not the test.
3.2.2 Variable’s influence on the PCE
In light of the results obtained in the previous section, it is worthwhile to study some of the top-performing ML models that characterize the behavior of OSCs. In this regard, standardized PFI and SHAP analyses are presented below, by means of Figure 9 and Figure 10, respectively. In this case, the behavior of OSCs is analyzed upon the variables of the dataset, for a temporal scope of up to 30 and 180 days, respectively.
Firstly, Figure 9 and Figure 9 compare the importance of all the variables without applying PFI filtering, for the best model at 30 and 180 days, in turn. It is confirmed that in initial stages, the effect of time is not clearly relevant on the PCE, compared to the model trained with data of 180 days. The remaining variables have similar implications, being the amount of solvent in the HTL layer, i.e., the amount of PEDOT:PSS, the more significant. Likewise, the P3HT:PCBM ratio also demonstrates certain relevance.
After applying PFI filtering, that is, removing variables with an effect on the R2 lesser than 4%, it is observed that in models generated with data up to 30 days, in Figure 9, the most significant variable is now the P3HT:PCBM ratio, above the amount of solvent in the HTL layer. Meanwhile, in models generated with data up to 180 days, in Figure 9, the effect of this quantity is greater. Additionally, dependencies with the value of PCBM are also observed, as it has more relevance for calculating the P3HT:PCBM ratio, given its non-linear influence since it modulates the value from the denominator.
Figure 9: PFI comparison between the best Machine Learning models generated with data up to: 30 days, in (a) and (c); 180 days, in (b) and (d). Variables with less than 4% of contribution to R2 are filtered out in (c) and (d).
Figure 10: SHAP comparison between the best Machine Learning models generated with data up to: 30 days, in (a) and (c); 180 days, in (b) and (d). Variables with less than 4% of contribution to R2 are filtered out in (c) and (d).
Regarding the SHAP analysis, it proves the extension of the insights previously extracted. When PFI filtering is not applied, Figure 10 and Figure 10 demonstrate the influence on the model of the amount of solvent PEDOT:PSS in the HTL layer, regardless of whether the model is generated with data up to 30 days or 180 days. However, in the latter case, the P3HT:PCBM ratio also shows greater relevance. As for the environmental conditions (temperature, humidity, dew point, and pressure), it is shown that they do not have wide influence. At initial stages (up to 30 days), there is low variation in these climate conditions, so that further insights should not be trustworthy made. By contrast, when enough time is taken into consideration (up to 180 days), it is confirmed that low humidity proves to be beneficial for these OSCs, which in our geographical location is normally correlated with high temperatures and atmospheric pressures. In a similar manner affects the dew point, which directly correlates with humidity. Nonetheless, the device encapsulation demonstrates that the effect of these variables is minimized, as it will be verified in the PFI analysis.
Finally, after applying PFI filtering, Figure 10 and Figure 10 highlight the same dependencies, where the effect of the PCBM value on the P3HT:PCBM ratio is again confirmed. As recently mentioned, it is verified that PFI filters out the effect of climate variables during the data acquisition, since it is mitigated by the encapsulation of the devices.
4 Discussion and conclusions
This paper has presented the application of an optimal ML framework to characterize the degradation behavior, in terms of PCE, of OSCs with multilayer structure ITO/PEDOT:PSS/P3HT:PCBM/Al. A dataset with electrical parameters of these devices has been acquired for more than 180 days, so as to come up with 996 entries, with variability in 7 variables. Through hyper-optimization of a set of well-recognized ML models we have presented an accuracy analysis of different methods, which were fed with OSC data periodized into sets from 30 to 180 days. The benchmarking has confirmed the validity of models like RF or GB to confer R2 values over 0.90, reaching in some extents R2$\sim$0.96-0.97. As for the error metrics obtained (RMSE, SSE and MAE), they are trustworthy bounded to $\lesssim$1of the target value (the PCE) when long term data is used for training. To reinforce the suitability of these ML models, classical LMS regression fitting methods have been presented. They demonstrated inoperability to perform reliably with multivariate models, neither with long-term data, and even less to estimate future values of the target PCE.
By contrast, our ML models offer high feasibility to predict the behavior of unknown OSCs, with accuracy of R2=0.97. Finally, the standardized result report provided by the ML framework, permits obtaining further insights of the dependencies of the variables of the dataset on the degradative behavior the OSCs, as per their importance in the ML models. It has been confirmed the importance of the solvent in the HTL layer, i.e., the amount of PEDOT:PSS. This is explained as the multilayer structure of the OSCs needs a minimum value of PEDOT:PSS to ensure a layer that completely covers the substrate. Moreover, the ratio P3HT:PCBM also exhibits significant importance, being higher values representative of greater impact on the model. It is worth noticing that models with data up to 180 days provide more stable implications in this sense. Finally, it has been validated that variables such as temperature, humidity, dew point and pressure have lesser impact on the models, explained by the encapsulation made to the OSCs during their manufacturing. Overall, this study comes up with robust and reliably ML models that characterize these OSCs in a reproducible, transparent and standardized manner, which are made publicly available.

### v3

@@
-Comparing Hyper-optimized Machine Learning Models for Predicting Efficiency Degradation in Organic Solar Cells
+General Machine Learning Models for Interpreting and Predicting Efficiency Degradation in Organic Solar Cells
 Abstract
-This work presents a set of optimal machine learning (ML) models to represent the temporal degradation suffered by the power conversion efficiency (PCE) of polymeric organic solar cells (OSCs) with a multilayer structure ITO/PEDOT:PSS/P3HT:PCBM/
-Al.
-To that aim, we generated a database with 996 entries, which includes up to 7 variables regarding both the manufacturing process and environmental conditions for more than 180 days.
+Within the range of renewable energy sources, photovoltaic (PV) plays a crucial role in addressing the exponential growth in global demand during this decade.
+Organic solar cells (OSCs) emerge as a promising alternative to silicon-based PVs due to their low cost, lightweight, and sustainable production.
+Despite achieving power conversion efficiencies (PCEs) exceeding values of 20%, OSCs still face challenges in stability and efficiency.
+In this context, recent advances in manufacturing, artificial intelligence (AI)-driven and machine learning (ML) models offer opportunities to optimize and screen OSCs for greater sustainability and commercial viability, thus potentially reducing costs while ensuring stable and long term performance.
+This work presents a set of optimal ML models to represent the temporal degradation suffered by the PCE of polymeric OSCs with a multilayer structure ITO/PEDOT:PSS/P3HT:PCBM/Al.
+To that aim, we generated a database with 166 entries with measurements of 5 OSCs, which includes up to 7 variables regarding both the manufacturing process and environmental conditions for more than 180 days.
 Then, we relied on a software framework that brings together a conglomeration of automated ML protocols that execute sequentially against our database by simply command-line interface.
-This easily permits hyper-optimizing and randomizing seeds of the ML models through exhaustive benchmarking so that optimal models are obtained.
-The accuracy achieved reaches values of the coefficient determination (R2) widely exceeding 0.90, whereas the root mean squared error (RMSE), sum of squared error (SSE), and mean absolute error (MAE) $\lesssim$1of the target value, the PCE.
-Additionally, we contribute with validated models able to screen the behavior of OSCs never seen in the database.
-In that case, R2$\sim$0.96-0.97 and RMSEsimilar-$\sim$1%, thus confirming the reliability of the proposal to predict.
-For comparative purposes, classical Bayesian regression fitting based on non-linear mean squares (LMS) are also presented, which only perform sufficiently for univariate cases of single OSCs.
-Hence they fail to outperform the breadth of the capabilities shown by the ML models.
-Finally, thanks to the standardized results offered by the ML framework, we study the dependencies between the variables of the dataset and their implications for the optimal performance and stability of the OSCs.
-Reproducibility is ensured by a standardized report altogether with the dataset, which are publicly available at Github [^1] https://github.com/davidumh11.
+This easily permits hyper-optimizing the ML models through exhaustive benchmarking so that optimal models are obtained.
+The accuracy for predicting PCE over time reaches values of the coefficient determination (R2) widely exceeding 0.90, whereas the root mean squared error (RMSE), sum of squared error (SSE), and mean absolute error (MAE) are significantly low.
+Additionally, we assessed the predictive ability of the models using an unseen OSC as an external set.
+For comparative purposes, classical Bayesian regression fitting based on non-linear least squares (LS) are also presented, which only perform sufficiently for univariate cases of single OSCs.
 keywords:
-Organic solar cell, Degradation, P3HT:PCBM, PEDOT:PSS, Machine Learning, Regression.
-journal: Engineering Applications of Artificial Intelligence
+Organic solar cells, power efficiency degradation, multilayer structure, Machine Learning, Regression.
+journal: Expert Systems with Applications
 […]
@@
 […]
-[label3]organization=Institute of Chemical Synthesis and Homogeneous Catalysis (ISQCH), CSIC, University of Zaragoza,addressline=Pedro Cerbuna 12, city=Zaragoza, postcode=50009, country=Spain
+[label3]organization=Department of Inorganic Chemistry, Institute of Chemical Synthesis and Homogeneous Catalysis (ISQCH), CSIC, University of Zaragoza,addressline=Pedro Cerbuna 12, city=Zaragoza, postcode=50009, country=Spain
 1 Introduction
-Nowadays electricity consumption has evolved exponentially over the last three decades, increasing from a total estimate of 15107.8 TWh to 28661 TWh between 2000 and 2022, respectively Rit2023 .
+Electricity consumption has evolved exponentially over the last three decades, increasing from a total estimate of 15277 to 29479 TWh between 2000 and 2023, respectively Rit2023 .
 One of the immediate consequences has been reflected in the price of this energy.
 Simultaneously, in the technological realm, there has been a notable shift in electricity generation towards renewable sources.
-Particularly, the maturity of conventional photovoltaic (PV) cells and the emergence of new organic and hybrid materials have made possible to get significant improvements in PV generation and its power conversion efficiency (PCE) NREL .
-However, despite the fact that the global consumption of PV energy increases from 1.08 TWh to 1310.02 TWh between 2000 and 2022 Rit2023 , its generation still falls short of meeting current demand.
+Particularly, the maturity of conventional photovoltaic (PV) cells and the emergence of new organic and hybrid materials have made it possible to get significant improvements in PV generation and its power conversion efficiency (PCE) NREL .
+However, even though the global consumption of PV energy has increased from 1.08 to 1629.9 TWh between 2000 and 2023 Rit2023 , its generation still falls short of meeting current demands.
 Hence, there are many challenges and open lines for improving the generation and efficiency of PV cells.
-In this context, organic solar cells (OSCs) have emerged as a promising alternative to silicon-based solar cells since, amongst other benefits, their production involves low-temperature manufacturing methods Seo2015 , reducing their carbon footprint.
+In this context, organic solar cells (OSCs) have emerged as a promising alternative to silicon-based solar cells since, amongst other benefits, their production involves low-temperature manufacturing methods Seo2015 and reduced carbon footprint.
 Their appeal also lies in the ease of processing, low cost, flexibility, and lightweight Yeh2013 .
 Moreover, the evolution of their PCEs has been much superior to classical technologies and it has been on par with other considered emerging technologies, presenting efficiencies that exceed values of 20% veintePCE .
-However, the commercialization of OSCs still faces various challenges, including stability and efficiency.
-Nevertheless, certain cost studies PVCost1 ; PVCost2 indicate that once these obstacles are overcome, OSCs can be manufactured at a cost lower than 1 dollar per peak Watt Bra2005 .
-This fact suggests that the optimization of the manufacturing procedure of these organic devices would certainly lead to reinforce their benefits, but also, and most importantly, their sustainability and stability to operate long-term feasibly.
-All in all, the potential capabilities of artificial intelligence (AI) protocols, through particular machine learning (ML)-based models, envisages great achievements in this sense.
-The integration of ML before transferring to industrial implementation it is being extensively tested these days.
+Even though the commercialization of OSCs still faces various challenges, including stability and efficiency, some cost studies PVCost1 ; PVCost2 ; FER_Heliyon indicate that once these obstacles are overcome, OSCs can be manufactured at a cost lower than one dollar per peak Watt Bra2005 .
+The integration of machine learning (ML) before transferring to industrial implementation is being extensively tested these days.
 For instance, within the field of cheminformatics, it is of paramount importance in the design of new materials.
-Numerous studies have demonstrated the potential of deep learning algorithms to analyze complex chemical data, thereby accelerating the process of discovering, for example, drugs, as well as identifying promising compounds with improved properties Lus2013 .
-Some of the early studies Han2013 ; Rup2012 laid the foundations for the application of ML models in predicting molecular properties, highlighting their great strengths.
-More recently, the deployment of AI techniques with high-performance prediction methods has enabled the emergence of software platforms capable of examining extensive chemical databases to identify potential materials with significant properties Unt2014 ; Wall2015 .
-A relevant database, which moved forward from general chemical structures to the inclusion of PV features was the project presented in Joh2011 .
-This work allowed to obtain custom predictive models that characterize the electrical behavior of PV devices in terms of parameters such as voltage, current, conversion efficiency, short-circuit current, open-circuit voltage, etc.
+Numerous studies have demonstrated the potential of deep learning algorithms to analyze complex chemical data, thereby accelerating the process of discovering, for example, drugs Dalmau2024 , as well as identifying promising compounds with improved properties Lus2013 ; Sanosa2024 .
+Some early studies Han2013 ; Rup2012 laid the foundations for the application of ML models in predicting molecular properties, highlighting their great strengths.
+More recently, the deployment of artificial intelligence (AI) techniques with high-performance prediction methods has enabled the emergence of software platforms capable of examining extensive chemical databases to identify potential materials with significant properties Unt2014 ; Wall2015 .
+Furthermore, researchers moved forward from general chemical structures to the inclusion of PV features Joh2011 , predicting the electrical behavior of PV devices in terms of parameters such as voltage, current, conversion efficiency, short-circuit current, open-circuit voltage, etc.
 Eib2021 ; Ser2016 ; Tol2019 ; Lop2016 .
-Thus there is an endless list of mathematical models mostly supported by ML and deep learning-based engines that combine different techniques, such as regression and statistical inference Moo2022 ; Eib2021 ; Mal2021 ; Seo2019 ; Sun2019 to optimize different target features of optoelectronic devices Miy2021 ; Mah2022a ; Mah2022b .
-Similarly, the growth of proposals with eco-design characteristics and AI support for the production of sustainable and durable energy generating materials is positively evolving Vel2023 ; Ju2018 ; Mah2023 .
-More specifically, some studies have trusted on ML for predicting optimal installation of energy harvesting systems HeliyonReview1 ; PV_forecast_EAAI ; HeliyonReview2 and consequently on predicting energy generation HeliyonPV ; ML_predictUSA ; PV_forecast_EAAI .
-Similarly, fault detection of these systems were also anticipated thanks to automated models DefectVisual ; DefectEAAI .
-Others delved into the manufacturing variables of their specific structures ML_structure ; Power_predict ; HeliyonPerov ; ML_OSC .
+Currently, there is a wide array of ML and deep learning-based engines Moo2022 ; Eib2021 ; Mal2021 ; Seo2019 ; Sun2019 designed to optimize different target features of optoelectronic devices Miy2021 ; Mah2022a ; Mah2022b .
+Other authors concentrated on AI-supported models to propose long lifecycle materials for sustainable energy generation  Vel2023 ; Ju2018 ; Mah2023 .
+Some studies have used ML for predicting optimal installation of energy harvesting systems  HeliyonReview1 ; PV_forecast_EAAI ; HeliyonReview2 whereas others have focused on predictors of energy generation HeliyonPV ; ML_predictUSA ; PV_forecast_EAAI .
+Likewise, fault detection of PV systems has been also anticipated thanks to automated models  DefectVisual ; DefectEAAI .
+Others delved into the manufacturing variables of the specific PV structures ML_structure ; Power_predict ; HeliyonPerov ; ML_OSC .
 In the same way, there is incipient research on the analysis of stability and durability of these devices PerovBig ; Lifetime ; PEDOT_stability ; PerovML .
 In sight of all these developments, this work explores different ML models to encode the PCE performance of polymeric OSCs based on multilayer ITO/PEDOT:PSS/P3HT:PCBM/Al.
-This permits producing models that reliable learn from the PCE evolution over time.
-Furthermore, several dependencies on the variables involved in the manufacturing PhysicaB_Dopant_study , but also on the environmental conditions DegradationVariables ; ThermalAnalisis ; DegradationHumidity , are also learnt by these models.
-To that aim, we have used an hyper-optimizer software, ROBERT ROBERT , that permits benchmarking with multiple ML models and so that to produce extensive and standardized results, which are generated with optimal methods.
-Our dataset consists of PCE measurements of 45 OSC devices with variability in up to 7 variables, acquired for more than 180 days.
-This means a database with 996 entries.
-The accuracy metrics demonstrate the validity of these models to learn and represent the temporal behavior of our OSCs, with determination coefficient up to R2=0.96, and root mean squared error (RMSE), sum of squared error (SSE) and mean absolute error (MAE) bounded around 1% of the target value, that is, the PCE.
-Additionally, traditional Bayesian regression models sustained by LMS have been introduced in order to confirm the advantages of the ML-automated protocols against these classical approaches.
-While these conventional methodologies such as PVBayes ; PVStatistical ; PVStatistical2 may yield satisfactory results in this domain, our approach demonstrates that the potential of computational learning modeling offers a much broader and comprehensive solution compared to those based on classical statistical models, showcasing, among other benefits, increased robustness and reliability for:
-  • Capturing multivariate relationships between variables, encompassing both linear and non-linear relationships.
-  • Learning from an entire dataset, not solely from a specific OSC device as in classical regression.
+To that aim, we have used the ROBERT program ROBERT , which automates data curation, screening of ML models, assessment of predictive ability, and feature analysis (for reproducibility details, see C).
+Our dataset consists of PCE measurements acquired for 180 days, of 5 OSC devices, resulting in a database with 166 entries.
+Although these devices are encapsulated, the database contains up to seven descriptors, some related to environmental conditions (temperature, hummidity, dew point and pressure) to inspect possible dependencies on their future performance and some others related to their composition and manufacturing (quantities of: solvent in the HTL layer, i.e., PEDOT:PSS; P3HT, PCBM; and volume ratio of P3HT:PCBM).
+The accuracy metrics demonstrate the validity of these models to learn and represent the temporal behavior of our OSCs, with determination coefficient up to R2=0.96, and low root mean squared error (RMSE), sum of squared error (SSE) and mean absolute error (MAE).
+Additionally, traditional Bayesian regression models sustained by non-linear least squares have been introduced to confirm the advantages of the ML-automated protocols against these classical approaches.
+While these conventional methodologies PVBayes ; PVStatistical ; PVStatistical2 may yield satisfactory results in univariate regression domains, our approach demonstrates that computational learning modeling offers a much broader and comprehensive solution compared to those based on classical statistical models.
+ML models show increased robustness and reliability for:
+  • Generating multivariate models to capture PCE behavior over time while ensuring the interpretability of variable influences.
+  • Learning from an entire dataset of multiple OSCs, not solely from a specific OSC device as in classical regression.
   • Screening new OSC devices not used during the learning phase, thereby predicting its behavior, as an unknown device.
-  • Predicting optimal variables to fabricate the OSC that maximizes the PCE and/or its stability over time.
-In contrast ML models, typical statistical algorithms are unable to predict; instead, they utilize estimations to extrapolate target values to immediately subsequent time instances, which must also be exclusive to a particular OSC rather than to an entire database with variability across different variables.
+  • Detecting optimal variables to fabricate the OSCs with optimal PCE and/or its stability over time.
 In summary, the key contributions of this research are as follows:
   1. 1.
-    Assessment of optimal ML models that learn the PCE behavior over time for polymer-based OSC devices with ITO/PEDOT:PSS/P3HT:PCBM/Al structure, considering various fabrication variables and environmental measurement conditions.
+    Assessment of optimal ML models that learn the PCE behavior over time for polymer-based OSC devices with ITO/PEDOT:PSS/P3HT:PCBM/Al structures.
   2. 2.
@@
   3. 3.
-    Identification of an optimal ML model capable of accurately predicting the PCE behavior of unseen OSC devices, unknown during the training and validation phases.
+    Identification of an optimal ML model capable of predicting the PCE behavior of unseen OSC devices.
   4. 4.
-    Analysis of dependencies among the dataset variables to establish influence on the performance of the OSCs devices and their fabrication.
+    Feature analysis to establish the influence of variables on the performance of the OSC devices.
   5. 5.
@@
 2.1 Organic Solar Cells
-The OSCs characterized in this work are manufactured through the spin-coating technique, which involves the deposition of overlapped polymeric thin films.
+The OSCs characterized in this work are manufactured through spin-coating technique, which involves the deposition of overlapped polymeric thin films.
 Polymers are deposited in solution, then a process of rotation followed by heating aids in removing the solvent, hence forming the layer.
@@
 As a substrate, a glass with a 60 nm thick semi-transparent ITO layer was used.
-The substrates are placed in the spinner with the ITO layer facing upwards so that the thin layers can be deposited on it.
-The first layer deposited was the PEDOT:PSS film.
-PEDOT:PSS was deposited at room temperature at a rotation speed of 6000 rpm for 60 seconds.
-The remaining solvent was removed by heating at 150C for ten minutes.
+The substrates were placed in the spinner with the ITO layer facing upwards so that the thin layers could be deposited on it.
+The first layer was the PEDOT:PSS film.
+It was deposited at room temperature at a rotation speed of 6000 rpm for 60 seconds.
+The remaining solvent was removed by heating at 150 C for 10 minutes.
 Once the first layer was deposited, the active layer composed by the P3HT:PCBM polymer blend was applied.
-This was done at 300 rpm for three minutes.
-It was then dried at 80C for one hour.
-The final layer deposited on the devices is the aluminium film.
-Aluminium is not deposited using the spin-coating technique but rather by metal evaporation in a vacuum chamber.
+This was done at 300 rpm for 3 minutes.
+It was then dried at 80 C for 1 hour.
+The final layer deposited on the devices was an aluminium film.
+Aluminium was not deposited using the spin-coating technique but rather by metal evaporation in a vacuum chamber.
 The equipment consists of a high vacuum chamber, two vacuum pumps, one of which is a high vacuum pump, and a power supply to provide the necessary current to evaporate the aluminium.
-Once the samples are placed inside the chamber, vacuum is achieved, reaching pressures of 10-6 mbar.
-Once this pressure is reached, aluminium is evaporated by Joule effect.
+Once the samples were placed inside the chamber, vacuum was achieved, reaching pressures of 10-6 mbar.
+When this pressure was reached, the aluminium evaporated by Joule effect.
 2.2 OSC database
-As initially mentioned in Section 1, our dataset consists of PCE measurements of 45 OSC devices where up to 7 variables are registered for more than 180 days.
-Electrical parameters (J-V curves) of these OSCs are acquired along with climate conditions.
+As initially mentioned in Section 1, our dataset consists of PCE measurements of 5 OSC devices, where up to 7 variables were registered for more than 180 days.
+Electrical parameters (J-V curves) of these OSCs were acquired along with climate conditions.
 The acquisition process comprises a Keithley-2400 equipment that acts as source generator for the voltage sweep as well as recorder of the generated PV current.
-The J-V curves have to be measured under light conditions (100 mW/cm2, AM 1.5G, and 25C) generated by a solar simulator Newport xenon arc lamp and an AM 1.5G filter.
-The electrical characterization is completed by determining the characteristic electrical parameters, including the short-circuit current density (Jsc), the open-circuit voltage (Voc), the maximum power point (Pmpp), the fill factor (FF), and the PCE.
-Obtaining these parameters allows to extract the PCE value as follows:
+The J-V curves were measured under light conditions (100 mW/cm2, AM 1.5 G, and 25 C) generated by a solar simulator Newport xenon arc lamp and an AM 1.5G filter.
+The electrical characterization was completed by determining the characteristic electrical parameters, including the short-circuit current density (Jsc), the open-circuit voltage (Voc), the maximum power point (Pmpp), the fill factor (FF), and the PCE.
+Obtaining these parameters allowed to extract the PCE value as follows:
 $PCE=\frac{P_{mpp}}{P_{inc}}=\frac{J_{sc}V_{oc}FF}{G}$ (1)
 being Pinc the incident solar power on the OSCs, which derives from the incident irradiance of the solar simulator, G.
-The block diagram of the equipment for the acquisition system is presented by Figure 1. This permitted obtaining a database with 996 entries. The variables and their units are presented by Table 1.
-Figure 1: Block diagram of the acquisition system used to generate the OSC database.
-Table 1: Variable details of the dataset.
+The block diagram of the equipment for the acquisition system is presented in Figure 1.
+This permitted obtaining a database with 166 entries.
+The variables and their units are presented in Table 1.
+Notice that, in order to get further insights on the future performance of the OSCs, environmental conditions at laboratory have also been measured, despite the devices were encapsulated.
+Figure 1: Block diagram of the acquisition system used to acquire the OSC database.
+Table 1: Detail of the dataset: Manufacturing variables and environmental conditions at laboratory.
 Variables Values
-Cell name [1-5]
 solvent quantity HTL (PEDOT:PSS) [${\mu}$l] [250-1000]
@@
 Time [days] [0-181]
-As an example, Figure 2 presents an OSC contained in the database, for which its current density (J) data have been periodically acquired against voltage (V) over 78 days.
-Specifically, Figure 2 displays the voltage range on the X-axis from 0 to 0.5 V, while Figure 2 adjusts the scale to show the range from 0 to 0.25 V, allowing for a clearer observation of the evolution of the J-V curve as the days pass by.
-The degradation principles of these organic devices suggest that over time, these curves, from which maximum power and energy conversion efficiency are obtained, should consistently decline.
-However, it can be observed in Figure 2 that there are certain temporal points between days 20, 26, 47 and 69, where the current density associated with specific voltages remains close to, or even exceeds, the current density associated with those same voltages but at earlier temporal points.
-This outline confirms that rather than time, there are other variables influencing the degradation of these devices, which implies that a specific OSC could exhibit slightly better performance in the future under certain conditions.
+As an example, Figure 2 presents an OSC device contained in the database, for which its current density (J) data have been periodically acquired against voltage (V) over time.
+Specifically, Figure 2 displays the voltage range on the X-axis from 0 to 0.5 V, while Figure 2 adjusts the scale to show the range from 0 to 0.25 V, allowing for a clearer observation of the evolution of the J-V curve when days go by (different colors of the legend).
+The degradation principles of these organic devices suggest that over time, these curves, from which maximum power and energy conversion efficiency are obtained, should consistently decay.
 Figure 2: Evolution of current density (J) vs. voltage (V) over time (days) for the OSC $Cell2$. (a) Curve J-V with V $\in$[0-0.7] V. (b) Same curve J-V with V $\in$[0-0.25] V.
-In the same manner, Figure 3 compares the temporal evolution of three different OSCs in terms of their normalized PCE values, ranging from 0 to 1, for more than 180 days.
+Likewise, Figure 3 compares the temporal evolution of three different OSCs in terms of their normalized PCE values, ranging from 0 to 1, for more than 180 days.
+Please note that normalization is used for visualization purposes, as typically used in this field Wurfel when analyzing PCE decay.
 Once again, it is confirmed that time leads to the degradation of the device’s conversion efficiency.
-However, it also emerges that differences in the fabrication of the various OSCs, as well as the environmental conditions during their measuring, may influence the trend followed by the PCE over time.
+However, it also emerges that differences in the fabrication of the OSCs, as well as the environmental conditions during their measuring, may influence the trend followed by the PCE over time.
 Figure 3: Evolution of the normalized PCE over time for three different OSCs.
-2.3 Regression Problem
-As already stated, this study aims to automatically model the degradative behavior exhibited by manufactured OSCs using ML approaches.
-It is clear that we deal with a regression problem, for which recent literature has demonstrated that is reliably and efficiently solved by computational learning.
-However, in order to establish further comparisons that highlight and reinforce the abilities of AI-based protocols, a preliminary step is carried out to present regression fitting using traditional LMS-supported Bayesian methods.
+2.3 Regression Problem with Classical Approaches
+This study aims to automatically model the degradative behavior exhibited by manufactured OSCs using ML approaches, which is a regression problem.
+In order to establish further comparisons that highlight and reinforce the abilities of ML-based protocols, a preliminary step was carried out to present regression fitting using traditional LS-supported Bayesian methods.
 Regression is a method for estimating the relationship between a response or output variable and one or more predictor or input variables.
@@
 From that starting point, a regression model relates response data to predictor data through one or more coefficients.
-Then a fitting algorithm is needed to calculate some model coefficients from a set of input data.
-In this sense, parametric fitting involves finding those coefficients, assuming that the data are statistical composed by deterministic and random components.
-Therefore, the parametric model estimates the deterministic component, whereas the random component is typically described as the error.
+Then a parametric fitting algorithm is needed to calculate some model coefficients from a set of input data.
+Therefore, the parametric algorithm estimates deterministic components, whereas random components are typically described as the error.
 Considering the model as a coefficient function of the independent variable, the error encodes random fluctuations around a Gaussian probability distribution.
-In many scopes, the goal is to minimize that error, classically addressed by means of LMS fitting approaches, such as: linear least squares, weighted least squares, robust least squares, non-linear least squares, etc.
+In many scopes, the goal is to minimize that error, classically addressed by means of least squares (LS) fitting approaches, such as: linear least mean squares, weighted least squares, robust least squares, non-linear least squares, etc.
 $y=f(X,\beta)+\varepsilon$ (2)
@@
 Whenever the fitting meets the specified convergence criteria, the final solution is assumed as valid.
-After observing in Section 2.2 the behavior of PCE over time for the various OSCs contained in our database, it was validated that the LMS regression fitting models yielding the best results were those with non-linear characterization.
+After observing in Section 2.2 the behavior of PCE over time for the various OSCs contained in our database, it suggests that the LS regression fitting models yielding the best results are those with non-linear characterization.
 Table 2 displays the selected models along with their expressions as a function of time, dependent on the adjustment coefficients.
 It should be noted that the capability of these classical models lies solely in modeling the univariate behavior of the time effect on the PCE values.
-While it is possible to explore other LMS regression fitting in the multivariate domain, they only allow for establishing linear relationships, which do not adequately capture the behavior of our devices.
-Table 2: LMS Bayesian regression fitting models to estimate PCE, denoted as $f(x)$, where $x$ represents time by means of parametric models.
+While it is possible to explore other LS regression fitting in the multivariate domain, they only allow for establishing linear relationships, which do not adequately capture the behavior of our devices.
+Table 2: LS Bayesian regression fitting models to estimate PCE, denoted as $f(x)$, where $x$ represents time by means of parametric models.
 Parametric model Coefficient expression of $f(x)$
@@
 $poly3$ $p_{1}x^{3}+p_{2}x^{2}+p_{3}x+p_{4}$
-2.4 Machine Learning framework
-In contrast to the previous classical approaches, ML moves forward to produce non-parametric regression models that adjust more complex behaviors without specifying the relationship between the output and input data by certain predetermined regression function.
-They may simply predict responses with new test data using a trained model.
-In this work, we exploit the advantages of a software framework, ROBERT ROBERT , developed under Python that facilitates hyper-optimization and benchmarking over well-recognized ML regression models by single command line instruction.
-This framework consists of the following modules:
-  • Data curation: It processes the target dataset in order to filter correlated descriptors, noise, duplicates, as well as to identify and to convert categorical variables into one-hot descriptors.
-  • Model selection: It iterates through multiple hyper-optimized models from scikit-learn pedregosa11a .
-It includes training size analysis by varying data partitions percentages (training-validation) and it also examines different data splitting options like k-neighbors clustering-based (KN) and random (RND).
-Permutation feature importance (PFI) detection is also performed.
-In consequence, models are computed with PFI-filtered descriptors and without them.
-As a default setting, this threshold eliminates features whose contribution to the model’s R2 is less than 4%.
-Moreover, it provides with insights on the dependencies of the input variables, thanks to SHapley Additive exPlanations (SHAP) analysis.
-Besides this, heatmap graphics are presented in order to compare models in terms of RMSE error.
-Amongst the generated models, it assesses the performance of Random Forests (RF), Multivariate Linear Models (MVL), Gradient Boosting (GB), Gaussian Process (GP), AdaBoost Regressor (AdaB), MLP Regressor Neural Network (NN), and Voting Regressor (VR) pedregosa11a .
-  • Prediction: Once the models are trained and validated, the framework permits external test sets to predict target values with the best models.
-  • Verify module: It assesses the predictive ability of the models, considering tests such as y-shuffle, y-mean, k-fold cross-validation, and one-hot features prediction.
-  • Report module: With the aim of providing reproducibility and transparency, this module offers a detailed report containing comprehensive information about the ML models utilized and replication instructions through command line executions.
-As an example of performance, Figure 4 illustrates the result obtained by ROBERT with our dataset as input data.
-Specifically, Figure 4 and Figure 4 compare the evolution exhibited by the ML models over time.
-The first figure depicts the result of a single model for training data up to 30 days, while the second figure does the same for training data up to 180 days.
-Their validity can be checked by inspecting the predicted outcome against the target values used as training (blue) and validation (yellow) data.
-Figure 4: Results of the best ML models, tested with training (blue) and validation (yellow) target data up to: (a) 30 days and (b) 180 days.
+2.4 ML framework
+In contrast to the previous classical approaches, ML moves forward to produce non-parametric regression models that adjust more complex behaviors.
+In this work, we exploit the advantages of a software framework developed under Python, ROBERT ROBERT (see C for reproducibility details), that facilitates hyper-optimization and benchmarking over well-recognized ML regression models by single command line instruction.
+This automated framework consists of the following modules (for further details please check the online documentation Readthedocs ):
+  • CURATE (data curation): It processes the input dataset in order to filter correlated descriptors, noise, duplicates, as well as to identify and to convert categorical variables into one-hot descriptors.
+  • GENERATE (model selection): It iterates through multiple hyper-optimized models from scikit-learn pedregosa11a , including Random Forest (RF), Multivariate Linear Model (MVL), Gradient Boosting (GB), Gaussian Process (GP), AdaBoost Regressor (AdaB), MLP Regressor Neural Network (NN), and Voting Regressor (VR).
+The algorithms are combined with different training-validation split sizes, from 60-40% to 90-10% using random data splitting.
+For each combination of algorithm and training size, two models are generated: one with all the descriptors and another with only the most important variables detected by permutation feature importance (PFI) analysis.
+Among all the possibilities, the program selects two optimal models based on RMSE error.
+  • PREDICT (external predictions selection): the framework is able to predict new target values.
+Moreover, it provides feature importances using SHapley Additive exPlanations (SHAP) and Permutance Feature Importance (PFI) analysis, and outlier detection.
+  • VERIFY (assesing predictive ability): It assesses the predictive ability of the models, considering tests such as y-shuffle, y-mean, k-fold cross-validation, and prediction with one-hot encoding.
+  • REPORT (generation of PDF reports): With the aim to enhance reproducibility and transparency, this module offers a detailed report containing comprehensive information about the ML models utilized and replication instructions through command line executions.
 3 Results
-This section introduces the results obtained using both classical LMS regression fitting models and the optimal ML models as generated by the framework used in this work, to estimate the temporal behavior of our OSCs.
+This section introduces the results obtained using both classical LS regression fitting and ML models to estimate the temporal behavior of our OSCs.
 The selected error metrics for analysis are briefly presented below:
   • Coefficient of determination R2: It quantifies how well the independent variables explain the variability of the dependent variable.
-Higher values indicate that the model fits the data well and captures larger proportion of the variability in the dependent variable.
+Higher values indicate that the model fits the data well and captures a larger proportion of the variability in the dependent variable.
     $R^{2}=1-\frac{SSE}{SS_{tot}}$ (5)
+    where SSE was defined in (4) and SStot is the total sum of squares.
   • RMSE: It provides a measure of the average magnitude of the errors made by the model in its predictions. Minimizing this error is often a goal when training regression models in ML.
     $RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$ (6)
+    where $n$ is the number of samples, $y_{i}$ is the real observed value and $\hat{y}_{i}$ is the predicted value.
   • MAE: Another common objective error that measures the average magnitude of the errors between the predicted values and the actual values of the target variable.
 Unlike RMSE, which penalizes large errors more heavily, MAE treats all errors equally by taking the average of their absolute values.
-This makes MAE more robust to outliers in the data compared to RMSE.
 $MAE=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$ (7)
-Meanwhile, SSE was defined in (4) and SStot is the total sum of squares, whereas $n$ is the number of samples, $y_{i}$ is the real observed value and $\hat{y}_{i}$ is the predicted value.
-3.1 Classical LMS regression fitting results
-Hence, Figure 5 presents the accuracy metrics of the five Bayesian regression models introduced in Table 2 to estimate the performance of PCE over 180 days, for each of the OSCs available in the database (Table 1).
+3.1 Classical LS regression fitting results
+Figure 4 presents the accuracy metrics of the five Bayesian regression models introduced in Table 2 to estimate the performance of PCE over 180 days, for each of the OSCs available in the database (Table 1).
 It is worth noting that these results correspond to the mean values, along with the standard deviation for each OSC.
-Specifically, the values of R2 are shown in Figure 5, RMSE in Figure 5, SSE in Figure 5, and MAE in Figure 5.
-It can be observed that these models perform effectively when the temporal range of the fitting data is limited: as long as the time goes by, all models experience an increase in error, while R2 decreases.
+In general terms, it can be observed that these models tend to perform slightly better over short time periods compared to the longer experiments.
+The maximum errors for all temporal fittings are bounded to values $\sim$0.06.
 According to the results of Figure 4, it is observed that on average, the $gauss2$ model, consisting of two Gaussian terms, is the one that best estimates the performance of the OSCs in terms of PCE over time.
-However, it should be noted that for a 30 day fit, this model requires more points to achieve a valid R2.
-On the other hand, it should be pointed out that the maximum errors for all temporal fittings are bounded to values $\sim$0.06, fact that represents an error $>$1% of the the target value, the PCE.
-Notwithstanding that, it must be reiterated the limitation of these methods to characterize the behavior of all devices contained in our database in a global manner.
-These results reflect solely the mean performance of fitting for single OSCs, in turn.
-Figure 5: Accuracy metrics of the LMS-supported Bayesian regression fitting over time.
+However, it should be noted that for a 30-day fit, this model requires more points to achieve a valid R2.
+It is worth mentioning that these methods cannot characterize the behavior of all devices contained in our database in a global manner, and the results obtained reflect solely the mean performance of fitting for individual OSCs.
+Figure 4: Accuracy metrics of the LS-supported Bayesian regression fitting over time.
 Five different models are evaluated: $exp1$blacksquare, $exp2$blacksquare, $gauss1$blacksquare, $gauss2$blacksquare, $poly3$blacksquare.
@@
 (d) Mean absolute error, MAE.
-Figure 6: PCE estimation over time with the best LMS-supported Bayesian regression fitting model ($gauss2$).
+3.1.1 PCE prediction with LS
+Once preliminary fittings have been made with traditional LS regression methods, it has been found that the best fitting corresponds to a double-term Gaussian function.
+Now we analyze the robustness of this method to predict the behavior of PCE against the temporal variable.
+For this purpose, four regression fittings have been obtained with PCE data acquired up to: 30, 60, 90, and 120 days, respectively.
+Then the behavior of PCE at 180 days has been predicted for each one (Figure 5 and Table 3).
+It can be confirmed that only the fitting with data up to 120 days is able to predict reliably future PCE values.
+This result demonstrates that even the best parametric LS fitting regression model requires more than half of the temporal data to reliably model the PCE evolution at future time values.
+Figure 5: PCE predition over time with the best LS-supported Bayesian regression fitting model ($gauss2$).
 Four temporal datasets are used to compute the fittings: 30 days $\circ$-- ∘ -; 60 days $\circ$-- ∘ -; 90 days $\circ$-- ∘ - and 120 days $\circ$-- ∘ -.
-3.1.1 PCE fitting estimation
-Once preliminary fitting have been made with traditional LMS regression methods, it has been found that the best fitting corresponds to a double-term Gaussian function.
-Now we analyze the robustness of this method to estimate the behavior of PCE against the temporal variable.
-For this purpose, different estimations have been obtained, as shown in Figure 6.
-This figure presents estimated PCE data from a specific OSC device, namely $Cell4$ in our databased, which has been used so far in the all the previous examples throughout this work.
-Here, the parametric model is obtained with fitting data at: 30, 60, 90, and 120 days, respectively.
-Then the behavior of PCE at 180 days has been estimated, in order to be tested against the real data of the database.
-It can be observed that the estimated trend for fitting models with 30, 60, and 90 days deviate considerably from the real behavior (represented by the black line).
-It can be concluded that fitting data up to at least 120 days are required to make a robust fitting that achieves sufficient confidence in estimating future data at 180 days.
-This fact demonstrates that even the best parametric LMS fitting regression model requires more than half of the temporal data to reliably model the PCE evolution at future time values.
-In order to get further insights on the estimations plotted in Figure 6 more easily, Table 3 presents the specific details of these estimations according to the error metrics RMSE, SSE, and MAE.
-Upon inspection, the perception extracted from the previous figure is validated: the errors are unacceptable for any metric unless the fit is made with more than half of the data (90 days), and only if the subsequent PCE estimation is tested over time values immediately posterior to the temporal limit of the fitting.
-Table 3: Accuracy metrics of the best LMS-supported Bayesian regression fitting model ($gauss2$) to estimate PCE over time, presented in Fig 6.
-Fitting data Estimation RMSE SSE MAE
+Table 3: Accuracy metrics of the best LS-supported Bayesian regression fitting model ($gauss2$) to predict PCE over time, presented in Figure 5.
+Fitting data Prediction RMSE SSE MAE
 30 days 60 days 0.0516 0.0213 0.0277
@@
 180 days 0.0138 0.0038 0.0105
-3.2 Machine Learning regression results
-In this section, we conduct the generation of ML models to learn from the temporal behavior of PCE versus time.
-It is noteworthy that, unlike the previous LMS-based regression fittings presented in Section 3.1, these ML models allow us to characterize the performance of our devices in a multivariate manner, thereby encoding the variability presented in their manufacturing, as well as in the particular environmental conditions at each temporal measurement, as indicated in Table 2.
-Thanks to the use of the tool ROBERT ROBERT , we achieve an exploration of various combinations of ML algorithms and partition sizes.
-In order to maintain the same accuracy estimation policy used in Section 3.1, Figure 7 presents the same metrics as Figure 5.
-In this case, for each temporal training dataset (X-axis), the coefficient R2 is presented in Figure 7, RMSE in Figure 7, SSE in Figure 7, and MAE in Figure 7, for the best ML model in each case.
-It can be observed that, in general, all errors are bounded approximately within ranges $\sim$[0.02-0.03], for training data of 120 days onwards.
-This implies an error percentage lower than 1%, with respect to the mean value of PCE for these OSCs, which is highly satisfactory for modeling.
-It is also worth noting the good performance of the RF model, with training-validation partitions of 90-10%.
-So as to comparatively assess the remaining ML models, Tables 45678, and 9, respectively present detailed accuracy results for the other models (GB, NN, MVL, and RF).
-Results with different training-validation ratios are presented: from 90-10% to 60-40%; while training data range from 30 to 180 days.
-For each table, the robustest model is highlighted in bold, considering the overall performance across all presented metrics, with priority given to the R2 value.
-It is worth mentioning that, even though they are not directly comparable models, it is confirmed that the errors of the ML methods are clearly bounded, while the LMS-based fitting regression errors grow with time.
-Furthermore, ML models have the ability to globally model a multivariate database, in contrast to univariate Bayesian fitting, which is independent for each device.
-Figure 7: Accuracy of the best Machine Learning regression models over time: 30 days RF-60-10 $\blacksquare$; 60 days RF-90-10 $\blacksquare$; 90 days RF-90-10 $\blacksquare$; 120 days RF-90-10 $\blacksquare$; 150 days RF-90-10 $\blacksquare$ and 180 days RB-90-10 $\blacksquare$.
+3.2 ML regression results
+Table 4: Accuracy metrics of the ML regression models computed with training data up to 180 days.
+Training data ML model R2 RMSE SSE MAE
+180 days GB-90-10 0.96 0.0640 0.0595 0.0520
+GB-80-20 0.96 0.0720 0.0600 0.0560
+GB-70-30 0.94 0.0840 0.0672 0.0650
+GB-60-40 0.95 0.0700 0.0431 0.0500
+NN-90-10 0.66 0.1400 0.3732 0.1100
+NN-80-20 0.76 0.1100 0.1138 0.0670
+NN-70-30 0.60 0.1600 0.3555 0.1200
+NN-60-40 0.43 0.2500 0.7265 0.1800
+MVL-90-10 0.81 0.1000 0.2929 0.0810
+MVL-80-20 0.81 0.1000 0.2783 0.0800
+MVL-70-30 0.74 0.1400 0.2696 0.1000
+MVL-60-40 0.74 0.1300 0.2802 0.0920
+RF-90-10 0.96 0.0480 0.0246 0.0390
+RF-80-20 0.97 0.0520 0.0309 0.0370
+RF-70-30 0.94 0.0670 0.0398 0.0460
+RF-60-40 0.93 0.0850 0.0644 0.0580
+In this section, we generate ML models to study the temporal behavior of PCE.
+Notice that, unlike the previous LS-based regression fittings presented in Section 3.1, these ML models allow us to characterize the performance of all the OSC devices under a single model, while also using the multiple variables in the dataset.
+These include manufacturing parameters and environmental conditions at each measurement.
+To comparatively assess the ML models, Table 4 presents detailed accuracy results for the algorithms GB, NN, MVL, and RF, with training data up to 180 days.
+Please note that all these models have PFI enabled, so that the most important features are considered.
+For further details about these features see Section 3.2.2 and C.
+The nomenclature for the models is: ALGORITHM-TRAINING-VALID (e.g.
+RF-90-10).
+Results with different training-validation ratios are presented: from 90-10% to 60-40%.
+In this study, the optimal ML models used RF algorithms with a 90-10 training-validation partitioning.
+It can be observed that, in general, all MAE and RMSE values are bounded within ranges $\sim$[0.02-0.03], for training data of 120 days onwards (Figure 6).
+The most robust model is highlighted in bold, considering the overall performance across all presented metrics, with priority given to the R2 value.
+Please note that the supplementary material (A) contains these same metrics for further inspection of models obtained with other temporal ranges.
+It is worth noting that errors of the ML methods are clearly bounded, regardless they operate on a multivariate database, in contrast to univariate Bayesian fitting presented in Section 3.1.
+Figure 6: Accuracy of the best ML regression models over time: 30 days RF-60-40 $\blacksquare$; 60 days RF-90-10 $\blacksquare$; 90 days RF-90-10 $\blacksquare$; 120 days RF-90-10 $\blacksquare$; 150 days RF-90-10 $\blacksquare$ and 180 days RB-90-10 $\blacksquare$.
 (a) Coefficient of determination, R2.
@@
 (d) Mean absolute error, MAE.
-Table 4: Accuracy metrics of the Machine Learning regression models computed with training data up to 30 days.
+3.2.1 PCE prediction with ML
+After comparatively evaluating the performance of ML models, this section validates their ability to predict the temporal behavior of the PCE for an OSC device never seen by the models.
+Firstly, the models are trained and validated without data of $Cell4$.
+Standing out above the others are the GB-90-10 and RF-90-10 models.
+In line with previous results (Table 4), and thus considering its robustness to model the entire dataset, RF-90-10 has been selected to get predictive inferences up to 180 days for the unseen OSC $Cell4$.
+Figure 7 provides such results.
+In particular, Figure 7 presents the results of the model when the dataset is splitted into traning and validation, at 90%-10% (blue and yellow points), whereas Figure 7 shows the prediction for $Cell4$ when its data are used as external test.
+Next, Figure 8 presents validation tests.
+Figure 8 shows the RMSE error of the model compared to those of the tests: y-mean, y-shuffle and onehot.
+Figure 8 produces a 5-fold cross-validation test CV5F , with satisfactory results.
+In consequence, neither data leakage nor overfitting is evidenced.
+Figure 7: Results of the best ML model, RF-90-10: (a) training-validation without data of $Cell4$ (R2=0.96, MAE=0.05, RMSE=0.071); (b) prediction of $Cell4$ as external test (R2=0.88, MAE=0.015, RMSE=0.021).
+Figure 8: Validation of the best ML model, RF-90-10: (a) RMSE values for validation tests y-mean, y-shuffle and onehot; (b) 5-fold cross-validation test (R2=0.89, MAE=0.06, RMSE=0.09).
+Finally, Figure 9 compares the temporal evolution of the PCE for $Cell4$ with the predicted data obtained by the model.
+Figure 9: PCE prediction over time with the best ML regression model (RF-90-10, blue) for an OSC not seen during the training-validation process ($Cell4$, black).
+3.2.2 Feature analysis
+Considering the results obtained in the previous section, it is worthwhile to study the top-performing ML model that characterizes the behavior of OSCs (i.e.
+RF-90-10).
+ML models trained with a limited number of data points are unable to characterize and predict with sufficient accuracy.
+For this reason, we focused on analyzing the complete dataset, with measures up to 180 days.
+In this regard, PFI and SHAP analyses are presented below, by means of Figure 10 and Figure 11, respectively.
+PFI evaluates the significance of individual features in a ML model by measuring the increase in the model’s prediction error after permuting the values of a specific feature.
+The resulting increase in error indicates the dependency on that feature PFI1 ; PFI2 .
+Besides, SHAP permits interpreting individual predictions of ML models by computing the contribution of each feature to the model’s as an additive feature attribution method SHAP2017 .
+Then Figure 10 presents the influence of the most relevant feature for the model, after applying PFI filtering, that is, removing variables with low effect on the R2.
+It is confirmed that the amount of solvent in the HTL layer, i.e.
+the amount of PEDOT:PSS, has the most significant influence.
+Likewise, the P3HT:PCBM ratio also demonstrates certain relevance.
+Additionally, dependencies with the value of PCBM are also observed, since it constraints the P3HT:PCBM ratio, given its non-linear influence in the denominator.
+As for the environmental conditions, it is known that low humidity proves to be beneficial for these OSCs, which in our geographical location is normally correlated with high temperatures and atmospheric pressures.
+In a similar manner it acts the dew point, which directly correlates with humidity.
+Nonetheless, the device encapsulation demonstrates that the effect of these variables is minimized, as PFI filtering proves that they are not relevant enough to influence substantially the model.
+Figure 10: PFI results for the best ML model generated with data up to 180 days.
+Regarding the SHAP analysis, Figure 11 validates the previous insights: high values of the amount of solvent PEDOT:PSS in the HTL layer, high ratios of P3HT:PCBM are relevant to get stable PCE values.
+Besides this, the effect of the PCBM value on the P3HT:PCBM ratio is again confirmed.
+Overall, the importance of PEDOT:PSS has been demonstrated as it plays an essential role in the multilayer structure of the OSCs.
+It confirms its relevance as the second layer to cover sufficiently the substrate.
+Moreover, the ratio P3HT:PCBM also demonstrates its role in dealing as charge carrier in the HTL layer.
+Figure 11: SHAP results for the best ML model generated with data up to 180 days.
+4 Discussion and conclusions
+This paper has presented the application of optimal ML frameworks to characterize the degradation behavior, in terms of PCE, of OSCs with multilayer structure ITO/PEDOT:PSS/P3HT:PCBM/Al.
+A dataset with 166 entries entries was created, containing PCE values of various OSC devices measured over 180 days.
+This dataset was supplemented with seven variables describing environmental conditions of the experiments and manufacturing parameters of the devices.
+Through hyper-optimization of a set of ML models we have presented an accuracy analysis of different methods, which were fed with OSC data periodized into sets from 30 to 180 days.
+The benchmarking has confirmed the validity of models like RF or GB to confer R2 values over 0.90, reaching in some extents R2$\sim$0.96-0.97 and error metrics (RMSE, SSE and MAE) significantly low when long term data is used for training.
+To reinforce the suitability of these ML models, classical LS regression fitting methods have been compared.
+These proved not to be suitable for a multivariable dataset like ours, especially when dealing with long term data of the OSCs.
+Consequently, their ability to predict PCE values is highly unreliable.
+ML models proved to offer high feasibility to predict the behavior of unknown OSCs.
+Feature analysis suggests that the most influential variable is the solvent in the HTL layer, i.e., the amount of PEDOT:PSS.
+This is explained as the multilayer structure of the OSCs needs a minimum value of PEDOT:PSS to ensure a layer that completely covers the substrate.
+Moreover, the ratio P3HT:PCBM also exhibits significant importance, being higher values representative of greater impact on the model.
+Finally, it has been observed that variables such as temperature, humidity, dew point and pressure have lesser impact on the models, explained by the encapsulation made to the OSCs during their manufacturing.
+CRediT authorship contribution statement
+Conceptualization D.V.
+and J.V.A.-R.; Methodology D.V., J.V.A.-R., D.D.
+and F.R.-M.; Software D.V., J.V.A.-R., D.D.
+and M.F.; Data curation D.V.
+and M.F.; Formal analysis D.V., F.R.-M.; Funding acquisition D.V., J.C.F.
+and J.V.A.-R.; Investigation D.V., F.R.-M.
+and J.V.A.-R.; Resources F.R.-M.
+and J.C.F.; Supervision J.C.F.
+and J.V.A.-R.; Validation D.D.
+and M.F.; Visualization D.V.
+and M.F; Writing - Original draft D.V.
+and J.V.A.-R.; Writing - Review & Editing J.V.A.-R., D.D.
+and F.R.-M.
+Acknowledgements
+This work is part of the project CIAICO/2023/193 funded by Generalitat Valenciana. The authors acknowledge that support.
+J.V.A.-R.
+and D.D.
+acknowledge Gobierno de Aragón Fondo Social Europeo (Research Group E07_23R), the State Research Agency of Spain (MCIN/ AEI/ 10.13039/501100011033/ FEDER, UE) for financial support (PID2022-140159NA-I00) and the European Union’s Recovery and Resilience Facility-Next Generation in the framework of the General Invitation of the Spanish Government’s public business entity Red.es to participate in talent attraction and retention programmes within Investment 4 of Component 19 of the Recovery, Transformation and Resilience Plan (MOMENTUM, MMT24-ISQCH-01).
+Appendix A Extended accuracy metrics of ML models
+This section comprises extended accuracy metrics of the ML models used in this work, when the temporal scope of the training-validation data is varied: 30, 60, 90, 120 and 150 days, respectively.
+Table 5: Accuracy metrics of the ML regression models computed with training data up to 30 days.
 Training data ML model R2 RMSE SSE MAE
@@
 RF-60-40 0.89 0.120 0.0725 0.0950
-Table 5: Accuracy metrics of the Machine Learning regression models computed with training data up to 60 days.
+Table 6: Accuracy metrics of the ML regression models computed with training data up to 60 days.
 Training data ML model R2 RMSE SSE MAE
@@
 RF-60-40 0.81 0.1300 0.0578 0.1000
-Table 6: Accuracy metrics of the Machine Learning regression models computed with training data up to 90 days.
+Table 7: Accuracy metrics of the ML regression models computed with training data up to 90 days.
 Training data ML model R2 RMSE SSE MAE
@@
 RF-60-40 0.74 0.1400 0.0860 0.1100
-Table 7: Accuracy metrics of the Machine Learning regression models computed with training data up to 120 days.
+Table 8: Accuracy metrics of the ML models computed with training data up to 120 days.
 Training data ML model R2 RMSE SSE MAE
@@
 RF-60-40 0.84 0.1000 0.0574 0.0770
-Table 8: Accuracy metrics of the Machine Learning regression models computed with training data up to 150 days.
+Table 9: Accuracy metrics of the ML regression models computed with training data up to 150 days.
 Training data ML model R2 RMSE SSE MAE
@@
 RF-60-40 0.90 0.0930 0.1069 0.0943
-Table 9: Accuracy metrics of the Machine Learning regression models computed with training data up to 180 days.
-Training data ML model R2 RMSE SSE MAE
-180 days GB-90-10 0.96 0.0640 0.0595 0.0520
-GB-80-20 0.96 0.0720 0.0600 0.0560
-GB-70-30 0.94 0.0840 0.0672 0.0650
-GB-60-40 0.95 0.0700 0.0431 0.0500
-NN-90-10 0.66 0.1400 0.3732 0.1100
-NN-80-20 0.81 0.1200 0.1067 0.0710
-NN-70-30 0.87 0.0860 0.0824 0.0660
-NN-60-40 0.89 0.0990 0.1021 0.0800
-MVL-90-10 0.81 0.1000 0.2929 0.0810
-MVL-80-20 0.80 0.1100 0.2645 0.0890
-MVL-70-30 0.78 0.1200 0.2626 0.0870
-MVL-60-40 0.74 0.1300 0.2802 0.0920
-RF-90-10 0.96 0.0280 0.0246 0.0290
-RF-80-20 0.94 0.0690 0.0391 0.0510
-RF-70-30 0.94 0.0670 0.0398 0.0460
-RF-60-40 0.93 0.0850 0.0644 0.0580
-3.2.1 PCE prediction
-After comparatively evaluating the performance of ML models, this section validates their ability to predict the temporal behavior of the PCE for an OSC device never seen by the models.
-In other words, a predictive inference of PCE is obtained for data from a particular OSC that has been omitted in the training and validation phases.
-Firstly, the models are trained and validated without data of $Cell4$ (same example used so far), thus the accuracy metrics are presented in Table 10.
-Standing out above the others are the GB-90-10 and RF-90-10 models.
-Secondly, the four trained models presented in Table 10, are queried to get predictive inferences up to 180 days for the unseen OSC, i.e.
-$Cell4$.
-Table 11 summarizes these accuracy results for each of these inferences.
-Again, both in Table 10 (training-validation) and in Table 11 (predictive testing) the highlighted error metrics together with the R2 values higher than 0.90, are indicative of feasibility to achieve our goal.
-Finally, Figure 8 provides visual detail for the particular results obtained with one of the best models, GB-90-10, to predict PCE over 180 days.
-Since the figure only shows the behavior of PCE against the time variable, once again, it is worth emphasizing on the potential of these models to encode accurately the multivariate variations of our dataset for an unknown OSC device.
-Table 10: Accuracy metrics of the Machine Learning regression models without training nor validation data of the OSC $Cell4$, which is not seen by the models.
-ML model R2 RMSE SSE MAE
-GB-90-10 0.88 0.069 0.0906 0.0600
-NN-90-10 0.89 0.095 0.3005 0.0740
-MVL-90-10 0.80 0.091 0.7528 0.0790
-RF-90-10 0.96 0.071 0.1118 0.050
-Table 11: Accuracy metrics of the Machine Learning regression models used as predictors of the PCE for the unseen OSC, $Cell4$.
-ML model R2 RMSE SSE MAE
-GB-90-10 0.97 0.055 0.0914 0.0370
-NN-90-10 0.80 0.1000 0.1975 0.0800
-MVL-90-10 0.90 0.1500 0.4010 0.1500
-RF-90-10 0.94 0.0610 0.0674 0.0470
-Figure 8: PCE prediction over time with the best Machine Learning regression model (GB-90-10, Table 11) for an OSC not seen during the training neither the validation not the test.
-3.2.2 Variable’s influence on the PCE
-In light of the results obtained in the previous section, it is worthwhile to study some of the top-performing ML models that characterize the behavior of OSCs.
-In this regard, standardized PFI and SHAP analyses are presented below, by means of Figure 9 and Figure 10, respectively.
-In this case, the behavior of OSCs is analyzed upon the variables of the dataset, for a temporal scope of up to 30 and 180 days, respectively.
-Firstly, Figure 9 and Figure 9 compare the importance of all the variables without applying PFI filtering, for the best model at 30 and 180 days, in turn.
-It is confirmed that in initial stages, the effect of time is not clearly relevant on the PCE, compared to the model trained with data of 180 days.
-The remaining variables have similar implications, being the amount of solvent in the HTL layer, i.e., the amount of PEDOT:PSS, the more significant.
-Likewise, the P3HT:PCBM ratio also demonstrates certain relevance.
-After applying PFI filtering, that is, removing variables with an effect on the R2 lesser than 4%, it is observed that in models generated with data up to 30 days, in Figure 9, the most significant variable is now the P3HT:PCBM ratio, above the amount of solvent in the HTL layer.
-Meanwhile, in models generated with data up to 180 days, in Figure 9, the effect of this quantity is greater.
-Additionally, dependencies with the value of PCBM are also observed, as it has more relevance for calculating the P3HT:PCBM ratio, given its non-linear influence since it modulates the value from the denominator.
-Figure 9: PFI comparison between the best Machine Learning models generated with data up to: 30 days, in (a) and (c); 180 days, in (b) and (d).
-Variables with less than 4% of contribution to R2 are filtered out in (c) and (d).
-Figure 10: SHAP comparison between the best Machine Learning models generated with data up to: 30 days, in (a) and (c); 180 days, in (b) and (d).
-Variables with less than 4% of contribution to R2 are filtered out in (c) and (d).
-Regarding the SHAP analysis, it proves the extension of the insights previously extracted.
-When PFI filtering is not applied, Figure 10 and Figure 10 demonstrate the influence on the model of the amount of solvent PEDOT:PSS in the HTL layer, regardless of whether the model is generated with data up to 30 days or 180 days.
-However, in the latter case, the P3HT:PCBM ratio also shows greater relevance.
-As for the environmental conditions (temperature, humidity, dew point, and pressure), it is shown that they do not have wide influence.
-At initial stages (up to 30 days), there is low variation in these climate conditions, so that further insights should not be trustworthy made.
-By contrast, when enough time is taken into consideration (up to 180 days), it is confirmed that low humidity proves to be beneficial for these OSCs, which in our geographical location is normally correlated with high temperatures and atmospheric pressures.
-In a similar manner affects the dew point, which directly correlates with humidity.
-Nonetheless, the device encapsulation demonstrates that the effect of these variables is minimized, as it will be verified in the PFI analysis.
-Finally, after applying PFI filtering, Figure 10 and Figure 10 highlight the same dependencies, where the effect of the PCBM value on the P3HT:PCBM ratio is again confirmed.
-As recently mentioned, it is verified that PFI filters out the effect of climate variables during the data acquisition, since it is mitigated by the encapsulation of the devices.
-4 Discussion and conclusions
-This paper has presented the application of an optimal ML framework to characterize the degradation behavior, in terms of PCE, of OSCs with multilayer structure ITO/PEDOT:PSS/P3HT:PCBM/Al.
-A dataset with electrical parameters of these devices has been acquired for more than 180 days, so as to come up with 996 entries, with variability in 7 variables.
-Through hyper-optimization of a set of well-recognized ML models we have presented an accuracy analysis of different methods, which were fed with OSC data periodized into sets from 30 to 180 days.
-The benchmarking has confirmed the validity of models like RF or GB to confer R2 values over 0.90, reaching in some extents R2$\sim$0.96-0.97.
-As for the error metrics obtained (RMSE, SSE and MAE), they are trustworthy bounded to $\lesssim$1of the target value (the PCE) when long term data is used for training.
-To reinforce the suitability of these ML models, classical LMS regression fitting methods have been presented.
-They demonstrated inoperability to perform reliably with multivariate models, neither with long-term data, and even less to estimate future values of the target PCE.
-By contrast, our ML models offer high feasibility to predict the behavior of unknown OSCs, with accuracy of R2=0.97.
-Finally, the standardized result report provided by the ML framework, permits obtaining further insights of the dependencies of the variables of the dataset on the degradative behavior the OSCs, as per their importance in the ML models.
-It has been confirmed the importance of the solvent in the HTL layer, i.e., the amount of PEDOT:PSS.
-This is explained as the multilayer structure of the OSCs needs a minimum value of PEDOT:PSS to ensure a layer that completely covers the substrate.
-Moreover, the ratio P3HT:PCBM also exhibits significant importance, being higher values representative of greater impact on the model.
-It is worth noticing that models with data up to 180 days provide more stable implications in this sense.
-Finally, it has been validated that variables such as temperature, humidity, dew point and pressure have lesser impact on the models, explained by the encapsulation made to the OSCs during their manufacturing.
-Overall, this study comes up with robust and reliably ML models that characterize these OSCs in a reproducible, transparent and standardized manner, which are made publicly available.
+Appendix B OSC database
+Database in $csv$ file consisting of 166 entries which includes up to seven variables regarding both the manufacturing process and environmental conditions for more than 180 days (Table 1).
+PCE values of several polymeric OSCs with a multilayer structure ITO/PEDOT:PSS/P3HT:PCBM/Al were measured.
+In addition, to ensure reproducibility, another $csv$ file provides a dataset without data of $Cell4$.
+Finally, another $csv$ file can be used to test the predictive ability of the model to predict data of such cell (see Section 3.2.1.)
+Appendix C ROBERT report
+Detailed report obtained with ROBERT (v1.0.6) after modeling the entire database. Results and reproducibility details are contained in this file.
+Data availability
+The program ROBERT (v1.0.6) to model the database is publicly available at: https://github.com/jvalegre/robert/releases
+The datasets are available in B.